OpenStudy (anonymous):

solve on the interval [0,2Pi): 2sintheta + square root of 3 = 0

OpenStudy (anonymous):

\[2\sin \theta = - \sqrt{3}\] \[\sin \theta = - \frac{ \sqrt{3} }{ 2 }\] Much easier to solve now

OpenStudy (anonymous):

take sin^-1 of - rad (3) / 2 There are 2 ans 1. 4pi/3, 5pi/3

OpenStudy (anonymous):

it's 4 pi/3, the 1. is just there c:

OpenStudy (anonymous):

whats that then

OpenStudy (anonymous):

i mean theta

OpenStudy (anonymous):

\[\frac{ 4\pi }{ 3 }\] and \[\frac{ 5 \pi }{ 3 }\] and the steps to do it is mentioned above

OpenStudy (anonymous):

thanks very much

OpenStudy (anonymous):

@abdollar19 welcome @Bebong thanks c:

OpenStudy (anonymous):

where u from